OpenStudy (anonymous):

Solve each of the quadratic equations below and describe what the solution(s) represent to the graph of each. x2 – 36 = 0 and x2 = 8x – 12 Part 2: Using complete sentences, compare and contrast the graphs of y = x2 − 36 and y + x2 = 8x − 12.

jhonyy9 (jhonyy9):

x=+/- 6 and to second x2-8x+12=0 x_1,_2=(8+/- sqrt(64-48))/2 = (8+/- 4)/2 = 2 and 6

OpenStudy (anonymous):

x^2-36=0 use difference of squares so (x-6)(x+6)=0 then what values of x make the equation true? in this case x=6,-6 the graph crosses the x axis at these points, because for these xs y=0. x^2=8x-12 x^2-8x+12=0 quadratic equation: x=(8+/-sqrt(8^2-4*1*12))/2 x=(8+/-sqrt(16))/2 x=4+/-2 x=2,6 so the graph crosses the x axis at 2 and 6

OpenStudy (anonymous):

the graph opens up, the x^2 term is positive, so it has a minimum point. vertex (x coordinate)=-b/2a -4/2=-2 plug it back into the original equation (-2,-16) x intercepts= x=-6,2

OpenStudy (anonymous):

Create your own unique quadratic equation in the form y = ax2 + bx + c that opens the same direction and shares one of the x-intercepts of the graph of y = x2 + 4x − 12. Explain whether the graph has a maximum or minimum point. Find the vertex and x-intercepts of the graph.

OpenStudy (anonymous):

make sure a is positive, and pick numbers so that when y=0, x=-6 or 2 will make the equation true. the graph has a minimum point, because it opens upward, and use the formula -b/2a to find the x coordinate of the vertex

OpenStudy (anonymous):

Example Please?